The points(2,8) and (a,29) are on a line of solope 14. find the value of a
Answer by Alan3354(69443)

The points(2,8) and (a,29) are on a line of solope 14. find the value of a
--------------
slope m = diffy/diffx
14 = (29-8)/(a-2)
14a - 28 = 21
a = 49/14 = 3.5